# Silent Guardian: Protecting Text from Malicious Exploitation by Large Language Models

### v3

@@ Abstract
To bridge this gap, we introduce Silent Guardian (SG), a text protection mechanism against LLMs, which allows LLMs to refuse to generate response when receiving protected text, preventing the malicious use of text from the source.
Specifically, we first propose the concept of Truncation Protection Examples (TPE). By carefully modifying the text to be protected, TPE can induce LLMs to first sample the end token, thus directly terminating the interaction. In addition, to efficiently construct TPE in the discrete space of text data, we propose a novel optimization algorithm called Super Taliored Protection (STP), which is not only highly efficient but also maintains the semantic consistency of the text during the optimization process.
@@ I Introduction
Recent advances in large language models (LLMs) have led to impressive performance in a variety of downstream language tasks, such as holding natural conversation [1], text and code generation [2, 3], and reading comprehension [4]. By automating tedious tasks and readily providing comprehensive information, they are expected to increase the productivity of society dramatically. However, while bringing convenience to people, this powerful ability also creates new security and privacy risks. For example, malicious users can use LLMs to exploit internet text to engage in illegal activities. Recent research [5] has indicated that by leveraging individual statements from social media, LLMs such as GPT-4 can accurately infer personal information such as gender, income, and location. This exposes personal privacy to enormous risks. In addition, the emergence of LLMs has automated the process of plagiarists and specialized artificial intelligence article rotation tools already exist [6], which poses a major challenge to copyright protection. Additionally, the capability of LLMs to mass-produce targeted rumors based on source texts [7] also makes governance in cyberspace increasingly difficult. It is worth noting that when LLMs have the ability to actively retrieve Internet texts, e.g., Bing Chat, highly automated malicious behaviors under the instructions of malicious users become possible, and the above risks are further amplified.
Therefore, there is an urgent need for a text protection mechanism to address these new risks in the context of LLMs to prevent text from being exploited maliciously. Unfortunately, while several attempts have been made to protect the copyright of text content, their main focus has been on traceability of unauthorized distribution or access control for unauthorized users. Specifically, some work adopts watermarking mechanisms for copyright protection, such as embedding digital watermarks within documents [8, 9, 10, 11] or visually adding real-time watermarks [12]. However, text watermarks can be easily bypassed by using LLMs for article spinning, which is a method of creating what looks like new content from existing content. In addition, some efforts use access restrictions to limit document replication and dissemination [13, 14], but the application scenarios for this mechanism are very limited and difficult to apply to today’s Internet landscape, where protected text content needs to be publicly released for a certain period of time. Moreover, some work proposes the use of obscuring and entities destroying to achieve the protection of personal privacy in texts [15, 16, 17, 18]. Notably, these methods usually require a large amount of manual review, resulting in low efficiency, lack of scalability, and difficulty in coping with rapidly growing large-scale data.
In general, the advent of LLMs has significantly broadened attackers’ perspectives, leading to a multitude of approaches in privacy attacks, rendering previous protection efforts no longer effective. The powerful ability of LLMs to understand, analyze, and create human language brings new challenges to text protection.
@@
Figure 1: Scenario of Silent Guardian. The adversary, upon acquiring the target text, articulates their requirements by adding prompt to the original text, thereby leading the model to produce harmful results. In Silent Guardian mechanism, STP aims to fine-tune the original text to prevent LLMs from generating any response. This kind of text is called TPE in this paper. The black arrows depict the adversary’s workflow, exemplifying three types of malicious operations: malicious analysis, plagiarism, and rumor fabrication, corresponding to the selection of personal privacy, copyrighted works, and government information, respectively. The green arrows represent the protective process of SG.
@@
Aiming to bridge these gaps, we propose a novel text protection mechanism against LLMs called Silent Guardian (SG), which can convert a piece of original text to protected text. Generally, when protected text is fed into LLMs as part of a prompt by malicious users, it will silence the LLMs, i.e., prevents them from generating any response and simply terminates the current conversation. We refer to such protected text as Truncation Protection Examples (TPE). Figure 1 provides an illustration of SG.
@@
Intuition: We mainly focus on auto-regressive language models, e.g., the GPT series, which generate a probability distribution for the next token after receiving a prompt. Then following a sampling method to obtain a specific token, this token is appended to the prompt for the next round of generation. This process will be repeated until one round of sampling results is a specific type of token known as the “end token”. Furthermore, prior research [19, 20, 21, 22, 23] has shown that LLMs may inherit the vulnerability of language models to adversarial examples [24, 25, 26, 27, 28, 29, 30]. When well-designed input text, i.e., an adversarial example for LLM, is fed into the LLMs, they can be induced to generate target content.
@@ II Related Work and Preliminary
Figure 2: An example of constructing TPE using STP on Vicuna. The blue part in the image represents the original text, and the red part represents the result after STP. $</s>$ represents the end token. After the token replacements shown in the box, this text successfully led the model to select the end token in the first sampling round. It can be observed that in the TPE constructed by STP, the model autonomously selects replacements such as letter casing changes and morphologically similar symbols (‘?’ to ‘?’). “[Prefixes]” and “[Suffixes]” represent additional requests that malicious users might add.
In this section, we review previous studies on traditional text protection, adversarial examples, and adversarial prompt against LLMs. And then introduce some notations of LLM in this paper.
@@ II-A Traditional Text Protection
Previous work on text protection primarily focused on two aspects of documents: copyright and privacy. For copyright, prevalent document protection methods involve embedding watermarks within the document, which includes image-based watermarking methods [12], semantics-based embedding [10, 31], and structure-based watermarking [8]. Additionally, there are protection methods that restrict unauthorized copying [13] or encode content in Unicode to create variations in copied text [14]. Privacy protection methods can generally be divided into two steps: identifying privacy entities within text and subsequently protecting privacy through methods such as substitution or masking [15, 16, 17, 18].
@@ II-B Adversarial Examples
However, the optimization problem in Equation 1 has been widely applied to continuous data such as images and speech, it does not directly apply to text data because the data space $X$ is discrete and the distance metric $\rho$ is difficult to define for text data. To circumvent these two issues, several attack algorithms at the character level, word level, and sentence level have been proposed. Character-level methods [24, 25] typically adjust characters through operations such as insertion, deletion, and swapping. Word-level methods create adversarial examples through word replacement, insertion, or deletion, using techniques like synonym replacement [33, 26], replacement with words close in embedding space [27], or leveraging language models to find the best replacement [28]. Some methods also focus on inserting or deleting words to construct adversarial examples [29]. Sentence-level methods perform extensive modifications at the sentence level [30], which can effectively disrupt model outputs but are less covert.
@@ II-C Adversarial Prompt against LLMs
With pre-trained language models [37, 38] becoming mainstream, prompt engineering [39] has become increasingly popular in recent years. However, recent research shows that through carefully constructed adversarial prompts, language models, including LLMs can be induced to output specified content. To achieve this, Autoprompt [19], GCG [20], and UAT [21] perform a greedy search to optimize the combination of tokens. PEZ [23] directly optimizes from the initial text, while GBDA [22] considers the adversarial example’s stealthiness and fluency but requires the introduction of additional models for constraints. These works explore classification tasks such as sentiment analysis, and natural language inference, as well as generative tasks such as red team testing and target generation.
@@ II-D Notations
Given a token sequence $[x_{1},x_{2},...,x_{n}]\in\mathcal{V}^{n}$, where $\mathcal{V}=\{token_{1},token_{2},...,token_{V}\}$ represents the set composed of all tokens in the vocabulary. $V$ and $n$ denote the size of the model’s vocabulary and the length of the token sequence respectively.
@@
After providing the probability prediction as described above, LLM can determine the next token $x_{n+1}$ through different sampling methods, and then add this token to the original token sequence to obtain a new token sequence $[x_{1},x_{2},...x_{n},x_{n+1}]$. LLM will repeat this process until a special token, the end token, is sampled.
@@ III Threat Model
    The architectures and parameters of the target LLMs are accessible, e.g., the open source LLMs.
@@ IV Silent Guardian
Existing text protection work cannot effectively address the issue of malicious exploitation of text by LLMs. To cope with this scenario, we introduce Silent Guardian (SG), a text protection mechanism against LLMs. The workflow of SG is to fine-tune the text to be protected into Truncation Protection Eeample (TPE) to prevent malicious exploitation by LLMs. Therefore, in this section, we will first introduce TPE, and then propose a novel algorithm to efficiently construct TPE, called Super Tailored Protection (STP).
@@ IV-A Truncation Protection Example
For the input $\mathcal{P}$, let the answer that selected end token in the first round of sampling be $r_{e}=[\text{end token}]$, $R_{\text{remain}}=R_{P}-r_{e}$. Then we can define this loss function as:
@@
Figure 3: The overview of Super Tailored Protection. (1) Calculate gradient for one-hot vectors: Convert the text to be protected into a one-hot vector representation. Input this into the LLM, and utilize the probability distribution of the predicted N+1th token and the end token’s probability distribution to compute the loss function. Calculate the gradient and propagate it backward. (2) Find the best replacement: Initially, generate semantically similar candidate set for each token in the text to be protected using neighboring tokens from the embedding layer. Then, take the results from step 1 to construct the final replacement candidate set from the semantically similar candidate set. Lastly, randomly select and identify the best replacement as the starting text for the next iteration.
@@ IV-B3 Gradient backpropagation
       $\text{loss}=\mathcal{L}(\mathcal{P})$ for $i=1,...,len(\mathcal{P})$ do
             $h_{i}=-\nabla_{v_{i}}\text{loss}$ $N_{i}=N(\mathcal{P}_{i})$ $S_{i}=Top-k(N_{i})$
      $\text{len}_{\text{part}}=\frac{B}{len(\mathcal{P})}$ for $b=1,...,B$ do
             $\lceil i=\frac{b}{\text{len}_{\text{part}}}\rceil$ $\tilde{\mathcal{P}}^{(b)}=\mathcal{P}$ $\tilde{\mathcal{P}}_{i}^{(b)}=Uniform(S_{i})$
@@ IV-B6 Random replacement and greedy search
First, divide $I$ into $\text{len}(\mathcal{P})$ parts, $I_{1},I_{2},...,I_{\text{len}(\mathcal{P})}$, each part corresponding to one changed position $i$.
Second, Randomly select tokens from $S_{i}$ for these positions to perform random replacements, which reduce time consumption for long protected text. Specifically, for $\tilde{\mathcal{P}}\in I_{i}$, let
@@ V-A1 Dataset
In theory, STP does not impose specific requirements on the theme or content of the prompt itself. To comprehensively validate the effectiveness of the defense, we selected 80 prompts across 9 different categories from Vicuna official website [40], which include writing, roleplay, common-sense, fermi, counterfactual, coding, math, generic, and knowledge.
In addition, we selected a set of texts from the novel Warden with varying lengths to verify the effectiveness of the text protection method on different text lengths. Specifically, we chose 10 texts each with approximately 40, 80, and 120 tokens, totaling 30 prompts. We denote this data set as a Novel dataset.
@@ V-A4 Metrics
    PSR, Success Rate of Truncation Protection, represents the effectiveness of truncation protection. We observed very few instances where the first round didn’t sample the end token but still stopped quickly. Therefore, The probability of the model sampling an end token as the first token can effectively describe PSR. For a more intuitive understanding, we define PSR as the probability of randomly sampling the end token in the first round at T=0. i.e.,
@@
    where $z_{i}$ represents the value in the model’s logits layer corresponding to $token_{i}$. A higher PSR indicates a more successful protection mechanism.
@@ V-A5 Baseline
We used GBDA [22] and PEZ [23] methods as baselines. Equation 6 presents the optimization objective of STP. In the baseline, we employed both PEZ and GBDA to achieve this optimization objective. The optimized initial value is set to the text to be protected. Then, we computed the loss function in Equation 6 and utilized PEZ and GBDA algorithms individually to optimize the prompt. It’s important to note that while GBDA method offers constraints on the concealment of adversarial examples, it requires a specific inference model. Hence, in this paper, we didn’t impose concealment constraints on GBDA. Additionally, to prevent gradient explosions when using these two methods, we used the Adam optimizer with values of epsilon (eps) $51e-$ 5.
@@ V-B1 White Box
Text
Can you help me write a resignation letter to my current employer, while leaving on good terms and expressing gratitude for the opportunities provided?
ShouldYou help me write a resignation letter to my current employer, while leaving on good terms but expressing grat attitude for the opportunities provided
Draft an apology email to a customer who experienced a delay in their order, and provide reassurance that the issue has been resolved.
@@
Write a compelling product launch announcement email to inform our customers of our new software solution.
Write a comp sell product launch announcement email to inform our customers of our new software answer\$?
Write a captivating movie review for a recently released science fiction film, discussing its plot, characters, and special effects.
Write a captvivating movie evaluate for an recently released science fiction film, discussizing its plot, character, and special effects}\$.
The Effectiveness of End Protection for Texts of Different Lengths. With increasing prompt length, the search space for constructing TPE grows and the concealment is enhanced under the same number of replacements. To investigate this, we conducted 15 rounds of TPE construction on the prepared Novel dataset. Table III shows the effectiveness of STP for texts of different token lengths. As expected, under the same number of rounds, the final effectiveness of TPE remains largely consistent. There’s even an improvement observed on LLaMA and Guanaco. Simultaneously, the concealment of TPE increases with the length of the text. This result means that our method has a distinctive advantage in protecting long texts and is highly suitable for real-world applications.
@@
Real-World Scenarios. In real-world scenarios, adversaries often add prefixes and suffixes to texts, such as “Please summarize the following text: [exploited text]” or “[exploited text]. Please summarize the preceding topic.” These are important steps when using LLMs for content generation. Therefore, In this section, we conducted experiments on Vicuna, LLaMA and Guanaco by adding prefixes “ Summarize following text,” “Summarize the topic of following text,” “Please summarize the topic of the following text and rewrite,” and suffixes “Summarize the preceding text,” “Summarize the topic of the preceding text,” “Please summarize the topic of the preceding text and rewrite” to the TPE on the Novel dataset. These are denoted as $\rm prefix_{1}$, $\rm prefix_{2}$, $\rm prefix_{3}$, $\rm suffix_{1}$, $\rm suffix_{2}$, and $\rm suffix_{3}$, respectively.
The experimental results are in Table IV and Table V. The PSR and PSR* in the table respectively represent the protective effect of TPE before and after adding a prefix or suffix. Even though we did not specifically optimize the protection for these prefixes and suffixes, the PSR remains at a relatively high level. Furthermore, as the length of the text to be protected increases, the impact of adding prefixes and suffixes diminishes on the protective effect. Additionally, longer prefixes and suffixes have a relatively minor effect on the protective outcome. This suggests that STP exhibit superior performance in tasks involving long texts and complex prefixes/suffixes, laying the groundwork for their scalability.

### v4

@@ Abstract
To bridge this gap, we introduce Silent Guardian (SG), a text protection mechanism against LLMs, which allows LLMs to refuse to generate responses when receiving protected text, preventing the malicious use of text from the source.
Specifically, we first propose the concept of Truncation Protection Examples (TPE). By carefully modifying the text to be protected, TPE can induce LLMs to first sample the end token, thus directly terminating the interaction. In addition, to efficiently construct TPE in the discrete space of text data, we propose a novel optimization algorithm called Super Tailored Protection (STP), which is not only highly efficient but also maintains the semantic consistency of the text during the optimization process.
@@ I Introduction
Recent advances in large language models (LLMs) have led to impressive performance in a variety of downstream language tasks [1], such as holding natural conversation [2], text and code generation [3, 4], and reading comprehension [5]. By automating tedious tasks and readily providing comprehensive information, they are expected to increase the productivity of society dramatically. However, while bringing convenience to people, this powerful ability also creates new security and privacy risks. For example, malicious users can use LLMs to exploit internet text to engage in illegal activities. Recent research [6] has indicated that by leveraging individual statements from social media, LLMs such as GPT-4 can accurately infer personal information such as gender, income, and location. This exposes personal privacy to enormous risks. In addition, the emergence of LLMs has automated the process of plagiarists and specialized artificial intelligence article rotation tools already exist [7], which poses a major challenge to copyright protection. Additionally, the capability of LLMs to mass-produce targeted rumors based on source texts [8] also makes governance in cyberspace increasingly difficult. It is worth noting that when LLMs have the ability to actively retrieve Internet texts, e.g., Bing Chat, highly automated malicious behaviors under the instructions of malicious users become possible, and the above risks are further amplified.
Therefore, there is an urgent need for a text protection mechanism to address these new risks in the context of LLMs to prevent text from being exploited maliciously. However, previous text protection methods mainly focused on copyright and privacy protection, each with its own shortcomings.
For text copyright protection, typical methods involve embedding watermark information into the text. This includes traditional synonym substitution-based methods [9, 10] and recently developed generative model-based watermarking methods [11, 12]. However, watermarking methods, being a passive defense, can only respond after an attack has occurred. Instead, some methods encode the content into Unicode, making it impossible for third parties to replicate it [13]. However, methods that restrict document copying can be easily bypassed by Optical Character Recognition (OCR).
For text privacy protection, existing methods mainly ensure the privacy of user text by changing the computation method or environment, such as fully homomorphic encryption [14], trusted execution environments [15], secure multi-party computation [16], and differential privacy [17]. However, fully homomorphic encryption requires excessive computational overhead, trusted execution environments, and secure multi-party computation, limited by communication and physical hardware constraints. Differentiated privacy may sacrifice the utility of the data in some cases.
@@
Figure 1: Scenario of Silent Guardian. The adversary, upon acquiring the target text, articulates their requirements by adding a prompt to the original text, thereby leading the model to produce harmful results. In the Silent Guardian mechanism, STP aims to fine-tune the original text to prevent LLMs from generating any response. This kind of text is called TPE in this paper. The black arrows depict the adversary’s workflow, exemplifying three types of malicious operations: plagiarism, rumor fabrication, and malicious analysis, corresponding to copyrighted works, government information, and personal privacy, respectively. The green arrows represent the protective process of SG.
@@
Aiming to bridge these gaps, we propose a novel text protection mechanism against LLMs called Silent Guardian (SG), which can convert a piece of the original text to protected text. Generally, when protected text is fed into LLMs as part of a prompt by malicious users, it will silence the LLMs, i.e., prevent them from generating any response and simply terminate the current conversation. We refer to such protected text as Truncation Protection Examples (TPE). Figure 1 provides an illustration of SG.
@@
Intuition: We mainly focus on auto-regressive language models, e.g., the GPT series, which generate a probability distribution for the next token after receiving a prompt. Then, following a sampling method to obtain a specific token, this token is appended to the prompt for the next round of generation. This process will be repeated until one round of sampling results is a specific type of token known as the “end token”. Furthermore, prior research [18, 19, 20, 21, 22] has shown that LLMs may inherit the vulnerability of language models to adversarial examples [23, 24, 25, 26, 27, 28, 29]. When well-designed input text, i.e., an adversarial example for LLM, is fed into the LLMs, they can be induced to generate target content.
@@ II Related Work and Preliminary
Figure 2: An example of constructing TPE using STP on Vicuna. $<\text{/s}>$ represents the end token. After the token replacements shown in the box, this text successfully led the model to select the end token in the first sampling round. It can be observed that in the TPE constructed by STP, the model autonomously selects replacements such as letter casing changes and morphologically similar symbols (‘?’ to ‘?’). “[Prefix]” and “[Suffix]” represent additional requests that malicious users might add.
In this section, we will review previous studies on traditional text protection, adversarial examples, and adversarial prompt against LLMs. Then, we will introduce some notations of LLM in this paper.
@@ II-A Traditional Text Protection
Previous work on text protection mainly focused on two aspects: copyright protection and privacy protection of the text.
For text copyright protection, typical methods involve embedding watermark information into the text. This includes traditional synonym substitution-based methods [9, 10] and recently developed generative model-based watermarking methods [11, 12]. However, watermarking methods, being a passive defense, can only respond after an attack has occurred. Instead, some methods encode the content into Unicode, making it impossible for third parties to replicate it [13]. However, methods that restrict document copying can be easily bypassed by Optical Character Recognition (OCR). The SG method is also an active defense method, but it generates adversarial content to prevent malicious analysis, making it less susceptible to perturbation by OCR or other methods, thus providing better protection.
For text privacy protection, existing methods mainly ensure the privacy of user text by changing the computation method or environment, such as fully homomorphic encryption [14], trusted execution environments [15], secure multi-party computation [16], and differential privacy [17]. However, fully homomorphic encryption requires excessive computational overhead, trusted execution environments, and secure multi-party computation, limited by communication and physical hardware constraints. Differentiated privacy may sacrifice the utility of the data in some cases. In contrast, SG can be performed in offline scenarios without the need for real-time computation. The computational overhead is acceptable in practical scenarios, and the environment deployment is simple and user-friendly. Compared to differential privacy, it also offers higher protection of text quality.
@@ II-B Adversarial Examples
However, the optimization problem in Equation 1 has been widely applied to continuous data such as images and speech. It does not directly apply to text data because the data space $X$ is discrete, and the distance metric $\rho$ is difficult to define for text data. To circumvent these two issues, several attack algorithms at the character level, word level, and sentence level have been proposed. Character-level methods [23, 24] typically adjust characters through operations such as insertion, deletion, and swapping. Word-level methods create adversarial examples through word replacement, insertion, or deletion, using techniques like synonym replacement [31, 25], replacement with words close in embedding space [26], or leveraging language models to find the best replacement [27]. Some methods also focus on inserting or deleting words to construct adversarial examples [28]. Sentence-level methods perform extensive modifications at the sentence level [29], which can effectively disrupt model outputs but are less covert.
@@ II-C Adversarial Prompt against LLMs
With pre-trained language models [35, 36] becoming mainstream, prompt engineering [37] has become increasingly popular in recent years. However, recent research shows that through carefully constructed adversarial prompts, language models, including LLMs, can be induced to output specified content. To achieve this, Autoprompt [18], GCG [19], and UAT [20] perform a greedy search to optimize the combination of tokens. PEZ [22] directly optimizes from the initial text, while GBDA [21] considers the adversarial example’s stealthiness and fluency but requires the introduction of additional models for constraints. These works explore classification tasks such as sentiment analysis and natural language inference, as well as generative tasks such as red team testing and target generation.
@@ II-D Notations
Given a token sequence $[x_{1},x_{2},...,x_{n}]\in\mathcal{V}^{n}$, where $\mathcal{V}=\{token_{1},token_{2},...,token_{V}\}$ represents the set composed of all tokens in the vocabulary. $V$ and $n$ denote the size of the model’s vocabulary and the length of the token sequence, respectively.
@@
After providing the probability prediction as described above, LLM can determine the next token $x_{n+1}$ through different sampling methods and then add this token to the original token sequence to obtain a new token sequence $[x_{1},x_{2},...x_{n},x_{n+1}]$. LLM will repeat this process until a special token, the end token, is sampled.
@@ III Threat Model
    The architectures and parameters of the target LLMs are accessible, e.g., the open-source LLMs.
@@ IV Silent Guardian
Existing text protection work cannot effectively address the issue of malicious exploitation of text by LLMs. To cope with this scenario, we introduce Silent Guardian (SG), a text protection mechanism against LLMs. The workflow of SG is to fine-tune the text to be protected as a Truncation Protection Eeample (TPE) to prevent malicious exploitation by LLMs. Therefore, in this section, we will first introduce TPE and then propose a novel algorithm to efficiently construct TPE, called Super Tailored Protection (STP).
@@ IV-A Truncation Protection Example
For the input $\mathcal{P}$, let the answer that selected end token in the first round of sampling be $r_{e}=[\text{end token}]$, $R_{\text{remain}}=R_{P}-r_{e}$. Then, we can define this loss function as:
@@
Figure 3: The overview of Super Tailored Protection. (1) Calculate gradient for one-hot vectors: Convert the text to be protected into a one-hot vector representation. Input this into the LLM and utilize the probability distribution of the predicted N+1th token and the end token’s probability distribution to compute the loss function. Calculate the gradient and propagate it backward. (2) Find the best replacement: Initially, generate a semantically similar candidate set for each token in the text to be protected using neighboring tokens from the embedding layer. Then, take the results from step 1 to construct the final replacement candidate set from the semantically similar candidate set. Lastly, randomly select and identify the best replacement as the starting text for the next iteration.
@@ IV-B3 Gradient backpropagation
       $\text{loss}=\mathcal{L}(\mathcal{P})$
       for $i=1,...,len(\mathcal{P})$ do
             $h_{i}=-\nabla_{v_{i}}\text{loss}$
             $N_{i}=N(\mathcal{P}_{i})$
             $S_{i}=Top-k(N_{i})$
      $\text{len}_{\text{part}}=\frac{B}{len(\mathcal{P})}$
       for $b=1,...,B$ do
             if $B>\text{len}(\mathcal{P})$ then
                  $i=\lceil\frac{b}{\text{len}_{\text{part}}}\rceil$
             else
                   repeat
                         $i=\text{random}(1,len(\mathcal{P}))$
                  until $i\neq\text{previous }i$
             $\tilde{\mathcal{P}}^{(b)}=\mathcal{P}$
             $\tilde{\mathcal{P}}_{i}^{(b)}=Uniform(S_{i})$
@@ IV-B6 Random replacement and greedy search
First, if $batch\ size>\text{len}(\mathcal{P})$, divide $I$ into $\text{len}(\mathcal{P})$ parts, $I_{1},I_{2},\ldots,I_{\text{len}(\mathcal{P})}$, each part corresponding to one changed position $i$. If $batch\ size\leq\text{len}(\mathcal{P})$, randomly select a unique position $i$ for each $\tilde{\mathcal{P}}^{j}\in I$.
Second, Randomly select tokens from $S_{i}$ for these positions to perform random replacements, which reduces time consumption for long protected text. Specifically, for $\tilde{\mathcal{P}}\in I_{i}$, let
@@ V-A1 Dataset
In theory, STP does not impose specific requirements on the theme or content of the prompt itself. To comprehensively validate the effectiveness of the defense, we selected 80 prompts across 9 different categories from Vicuna official website [38], which include writing, roleplay, common-sense, fermi, counterfactual, coding, math, generic, and knowledge. We denote this data set as the Vicuna dataset.
In addition, we selected a set of texts from the novel Warden with varying lengths to verify the effectiveness of the text protection method on different text lengths. Specifically, we constructed eight sets of texts of different lengths, each of which is about 40, 80, 120, 160, 200, 240, 280, and 320 tokens long. Each group has 10 texts, totaling 80 texts. We denote this data set as the Novel dataset.
@@ V-A4 Metrics
    PSR, Success Rate of Truncation Protection, represents the effectiveness of truncation protection. We define PSR as the softmax score corresponding to the end token, i.e.,
@@
    where $z_{i}$ represents the value in the model’s logits layer corresponding to $token_{i}$. A higher PSR indicates a more successful protection effectiveness.
@@ V-A5 Baseline
We used GBDA [21] and PEZ [22] methods as baselines. Equation 6 presents the optimization objective of STP. For baseline, we employed both PEZ and GBDA to achieve this optimization objective. The optimized initial value is set to the text to be protected. Then, we computed the loss function in Equation 6 and utilized PEZ and GBDA algorithms individually to optimize the prompt. It’s important to note that while the GBDA method offers constraints on the concealment of adversarial examples, it requires a specific inference model. Hence, in this paper, we didn’t impose concealment constraints on GBDA. Additionally, to prevent gradient explosions when using these two methods, we used the Adam optimizer with values of epsilon (eps) $51e-$ 5.
@@ V-B1 White Box
Original Text Protected Text
Can you help me write a resignation letter to my current employer, while leaving on good terms and expressing gratitude for the opportunities provided? ShouldYou help me write a resignation letter to my current employer, while leaving on good terms but expressing grat attitude for the opportunities provided
@@
Write a compelling product launch announcement email to inform our customers of our new software solution. Write a comp sell product launch announcement email to inform our customers of our new software answer\$?
Write a captivating movie review for a recently released science fiction film, discussing its plot, characters, and special effects. Write a captvivating movie evaluate for an recently released science fiction film, discussizing its plot, character, and special effects}\$.
The Effectiveness of Truncation Protection for Texts of Different Lengths. With increasing prompt length, the search space for constructing TPE grows, and the concealment is enhanced under the same number of replacements. To investigate this, we conducted 15 rounds of TPE construction on the prepared Novel dataset. Table III shows the effectiveness of STP for texts of different token lengths. As expected, under the same number of rounds, the final effectiveness of TPE remains largely consistent. An improvement was even observed in LLaMA and Guanaco. Simultaneously, the concealment of TPE increases with the length of the text. This result means that our method has a distinctive advantage in protecting long texts and is highly suitable for real-world applications.
@@
Real-World Scenarios. In real-world scenarios, adversaries often add prefixes and suffixes to texts, such as “Please summarize the following text: [exploited text]” or “[exploited text]. Please summarize the preceding topic.” These are important steps when using LLMs for content generation. Therefore, In this section, we conducted experiments on Vicuna, LLaMA, and Guanaco by adding prefixes “Summarize following text:” “Summarize the topic of following text:” “Please summarize the topic of the following text and rewrite:” and suffixes “Summarize the preceding text.” “Summarize the topic of the preceding text.” “Please summarize the topic of the preceding text and rewrite.” to the TPE on the Novel dataset. These are denoted as $\rm prefix_{1}$, $\rm prefix_{2}$, $\rm prefix_{3}$, $\rm suffix_{1}$, $\rm suffix_{2}$, and $\rm suffix_{3}$, respectively.
The experimental results are in Table IV and Table V. The PSR and PSR* in the table, respectively, represent the Protection Success Rate of TPE before and after adding a prefix or suffix. Even though we did not specifically optimize the protection for these prefixes and suffixes, the PSR remains at a relatively high level. Furthermore, as the length of the text to be protected increases, the impact of adding prefixes and suffixes diminishes on the protective effect. Additionally, longer prefixes and suffixes have a relatively minor effect on the protective outcome. This suggests that STP exhibit superior performance in tasks involving long texts and complex prefixes/suffixes, laying the groundwork for their scalability.